The Holtzman Corporation has assets of $410,000, current liabilities of $69,000, and long-term liabilities of $89,000. There is $37,500 in preferred stock outstanding; 20,000 shares of common stock have been issued.  

a. Compute book value (net worth) per share. (Round your answer to 2 decimal places.)
  



b. If there is $33,600 in earnings available to common stockholders, and Holtzman’s stock has a P/E of 22 times earnings per share, what is the current price of the stock? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
  


c. What is the ratio of market value per share to book value per share? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
Answer #1
a.
Total stockholders' equity = Total assets - Current liabilities - Long term liabilities = 410000 - 69000 - 89000 252000
Book value per share = ( Total stockholders' equity - Preferred stock ) / Shares of common stock outstanding = ( 252000 - 37500 ) / 20000 10.73
b.
Earnings per share = Earnings available to common stockholders / Shares of common stock outstanding = 33600 / 20000 1.68
P/E ratio = Current stock price / Earnings per share
22 = Current stock price / 1.68
Current stock price = 22 * 1.68 36.96
c.
Market value to book value = Current stock price / Book value per share = 36.96 / 10.73 3.44 times